# Bidirectional Consistency Models

### v2

@@ 3.2 Bidirectional Consistency Training
Model Illustration of the Training Objective Detailed Form of Loss
Table 1: Comparison of consistency training and bidirectional consistency training. The figures illustrate the main objective of each method, where $\bar{\bm{\theta}}$ means stop gradient over ${\bm{\theta}}$. Note that for BCM, there are two possible scenarios corresponding to the denoising and diffusion direction, respectively.
We now discuss the training of BCM, which we dub as Bidirectional Consistency Training (BCT). Following Song et al. [17], Song and Dhariwal [18], we discretize the time horizon $[\varepsilon,T]$ into $N-1$ intervals, with boundaries $\varepsilon=t_{1}<t_{2}<\dots<t_{N}=T$, and increase $N$ gradually during training.
@@
where ${\mathbf{x}}$ is one training sample, ${\mathbf{z}}\sim\mathcal{N}(0,{\bm{I}})$, and $\bar{\bm{\theta}}$ is the stop gradient operation on $\bm{\theta}$, and $\lambda(t_{n})=\nicefrac{{1}}{{|t_{n}-t_{n+1}|}}$. Note, that we replace $\bm{\theta}^{-}$ in Equation 6 with $\bar{\bm{\theta}}$ according to Song and Dhariwal [18].
@@
[…]
@@
We denote this as a “soft” trajectory constraint. However, ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{u}$ generated by the diffusion SDE do not necessarily lie on the same PF ODE trajectory, and hence Equation 13 still fails to build the desired bidirectional consistency. Instead, we note that when the CT loss defined in Equation 11 converges, we have $\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}_{u},u,0)\approx\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}_{t},t,0)\approx{\mathbf{x}}$. We therefore optimize:
@@
Empirically, we found Equation 14 plays a crucial role in ensuring accurate inversion performance. We provide experimental evidence for this loss choice in Section D.1.
@@
$\begin{split}\mathcal{L}_{BCT}^{N}(\bm{\theta})=\mathbb{E}_{{\mathbf{z}},{\mathbf{x}},t_{n},t_{n^{\prime}}}\Big{[}\lambda(t_{n}&)d(\bm{f}_{\bm{\theta}}({\mathbf{x}}+{t_{n+1}}{\mathbf{z}},t_{n+1},0),\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}+{t_{n}}{\mathbf{z}},t_{n},0))\\ &+\lambda^{\prime}(t_{n},t_{n^{\prime}})d\left(\bm{f}_{\bar{\bm{\theta}}}(\bm{f}_{\bm{\theta}}({\mathbf{x}}+t_{n}{\mathbf{z}},t_{n},t_{n^{\prime}}),t_{n^{\prime}},0)),\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}+t_{n}{\mathbf{z}},t_{n},0)\right)\Big{]}\end{split}$ (15)
@@
Regarding other training settings, encompassing the scheduler function for $N$, the sampling probability for $t_{n}$ (aka the noise schedule $p(n)$ in [17, 18]), the EMA rate, and more, we follow Song and Dhariwal [18]. We summarize our training procedure in Algorithm 1 , with more details regarding the hyperparameters in Section A.1.
As a further illustration, we compare CT and BCT in Table 1. For completeness, we also compare them with CTM in Table 4 in Appendix C.
@@ 3.3 Sampling
In this section, we detail the sampling schemes of BCM. First, BCM supports 1-step sampling, similar to CMs and CTM. Additionally, BCM’s capability to navigate both forward and backward along the PF ODE trajectory allows us to design multi-step sampling strategies to improve the sample quality. We present two schemes and a combined approach that has empirically demonstrated superior performance in the following.
3.3.1 Ancestral Sampling
@@
The most straightforward way for multi-step sampling is to remove noise sequentially. Specifically, we first divide the time horizon $[0,T]$ into $N$ sub-intervals, marked by the boundaries $0=t_{0}<t_{1}<\dots<t_{N}=T$. It is worth noting that this discretization strategy may differ from the one used during the training of BCM. Then, we sample a noise image ${\mathbf{x}}_{T}\sim\mathcal{N}(0,T^{2}{\bm{I}})$, and sequentially remove noise with the network:
${\mathbf{x}}_{t_{n-1}}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t_{n}},t_{n},t_{n-1}),\quad n=N,N-1,...,1.$ (16)
Since this sampling procedure can be viewed as drawing samples from the conditional density $p_{{\mathbf{x}}_{t_{n-1}}|{\mathbf{x}}_{t_{n}}}({\mathbf{x}}_{t_{n-1}}|{\mathbf{x}}_{t_{n}})$, we dub it as ancestral sampling, and summarize it in Algorithm 3. Note, that we can view 1-step sampling as a special case of ancestral sampling where we only divide the time horizon into a single interval.
3.3.2 Zigzag Sampling
@@
Song et al. [17] proposed another effective sampling method (Algorithm 1 in [17]), iteratively re-adding noise after denoising, which we dub as zigzag sampling hereafter. Similar to ancestral sampling, we also define a sequence of time steps $t_{1}<\dots<t_{N}=T$. However, different from ancestral sampling where we gradually remove noise, we directly map ${\mathbf{x}}_{T}$ to ${\mathbf{x}}_{0}$ by $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{T},T,0)$. Subsequently, we add a fresh Gaussian noise to ${\mathbf{x}}_{0}$, mapping it from time 00 to time $t_{n-1}$, i.e., ${\mathbf{x}}_{t_{n-1}}={\mathbf{x}}_{0}+t_{n-1}\bm{\sigma}$, where $\bm{\sigma}\sim\mathcal{N}(0,{\bm{I}})$. This process repeats in this zigzag manner until all the designated time steps are covered.
According to Song and Dhariwal [18], the two-step zigzag sampler effectively reduces FID and is theoretically supported [25]. However, as Kim et al. [19] have highlighted, the process of injecting fresh noise can alter the content of the image after each iteration, which will be undesired, especially considering that our tasks will both include generation and inversion. One may immediately think that setting the injected noise to be the same as the initial random noise can fix this issue. However, we reveal that this will significantly damage the quality of the generated images.
@@
Fortunately, our proposed BCM provides a direct solution to this dilemma, leveraging its capability to traverse both forward and backward along the PF ODE. Rather than manually reintroducing a large amount of fresh noise, we initially apply a small amount and let the network amplify it. Given the small scale of the initial noise, this method maintains a nearly deterministic sampling process, thereby leveraging the benefits of both fresh noise and determinacy. In a nutshell, for iteration $n$ ($n=N,N-1,\cdots,2$), we have
${\mathbf{x}}_{0}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t_{n}},t_{n},0),\quad\bm{\sigma}\sim\mathcal{N}(0,{\bm{I}}),\quad{\mathbf{x}}_{\varepsilon_{n-1}}\leftarrow{\mathbf{x}}_{0}+\varepsilon_{n-1}\bm{\sigma},\quad{\mathbf{x}}_{t_{n-1}}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{\varepsilon_{n-1}},\varepsilon_{n-1},t_{n-1}).$ (17)
@@
where $\varepsilon_{n-1}$ is the scale of the small noises we add in $n$-th iteration. We detail this scheme in Algorithm 4. Also, to verify its effectiveness, in Figure 2, we illustrate some examples to compare the generated images by 1) manually adding fresh noise, 2) manually adding fixed noise, and 3) our proposed sampling process, i.e., adding a small noise and amplifying it with the network. We can clearly see our method maintains the generated content.
Algorithm 2 Combination of ancestral and zigzag sampling
Generated image ${\mathbf{x}}$.
${\mathbf{x}}_{t_{N}}\leftarrow{\mathbf{x}}_{T}$.
# Ancestral sampling steps
for $n=N,\cdots,2$ do
end for
# Zigzag sampling steps
${\mathbf{x}}_{\tau_{M}}\leftarrow{\mathbf{x}}_{t_{1}}$.
for $m=M,\cdots,2$ do
end for
${\mathbf{x}}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{\tau_{1}},\tau_{1},0)$.
Return: ${\mathbf{x}}$.
@@
Figure 2: Comparison of different strategies of adding fresh noise in zigzag sampling. (a) 1-step generation. (b) Zigzag sampling with manually added fresh noise. We can see that the new noises drastically alter the content. (c) Zigzag sampling with manually added noise. But to enforce determinacy, we fix the injected fresh noise in each iteration to be the same as the initial one. We can see that although we get the same content, the quality significantly deteriorates. (d) Zigzag sampling with BCM. At each iteration, we apply a small amount of noise and let the network amplify it. We can see that the image content is mostly maintained.
3.3.3 Combination of Both
@@
Although BCM’s zigzag sampling significantly enhances sample quality, Kim et al. [19] note that these long jumps along the PF ODE can lead to accumulative errors, especially at high noise levels. This, in turn, hampers further improvements in sample quality. Therefore, we propose a combination of ancestral sampling and zigzag sampling. Specifically, we first perform ancestral sampling to rapidly reduce the large initial noise to a more manageable noise scale and then apply zigzag sampling within this reduced noise level. We describe this combined process in Algorithm 2, and empirically find it results in superior sample quality compared to employing either ancestral sampling or zigzag sampling in isolation.
@@ 3.4 Inversion
BCM inverts an image following the same principle of sampling. Specifically, we also set an increasing sequence of noise scales $\varepsilon=t_{1}<t_{2}<\cdots<t_{N}\leq T$. Note that, in contrast to the generation process, it is not always necessary for $t_{N}$ to equal $T$. Instead, we can adjust it as a hyperparameter based on the specific tasks for which we employ inversion. Then, given an image ${\mathbf{x}}_{0}$, we first inject a small Gaussian noise by ${\mathbf{x}}_{t_{1}}={\mathbf{x}}_{0}+\varepsilon\bm{\sigma}$, and then sequentially add noise with the network, i.e.,
${\mathbf{x}}_{t_{n+1}}=f({\mathbf{x}}_{t_{n}},t_{n},t_{n+1}),\quad n=1,2,\cdots,N-1.$ (18)
The adoption of small initial noise is due to the observation that the endpoint of the time horizon is less effectively covered and learned during training, as discussed in Section D.3. Empirically, we find this minor noise does not change the image’s content and leads to lower reconstruction errors when $\varepsilon=0.07$. One may also include denoising steps interleaved with noise magnifying steps, like zigzag sampling, but we find it helps little in improving inversion quality. We summarize the inversion procedure in Algorithm 5.
@@ 4 Experiments and Results
In this section, we first present the results of the two basic functions of BCM, i.e., generation and inversion. In Section 4.3, we showcase some potential applications enabled by these two functions.
@@ 4.1 Image Generation
Diffusion with Fast Sampler / Distillation
DDIM [6] 10 8.23 -
DPM-solver-fast [26] 10 4.70 -
Progressive Distillation [27] 1 8.34 8.69
CD (LPIPS) [17] 1 3.55 9.48
@@
Score SDE (deep) [7] 2000 2.20 9.89
DDPM [5] 1000 3.17 9.46
EDM [15] 35 2.04 9.84
CT (LPIPS) [17] 1 8.70 8.49
@@
CTM (LPIPS, GAN loss) [19] 1 2.39 -
iCT [18] 1 2.83 9.54
@@
Table 2: NFE and sample quality on CIFAR-10. Results estimated from Figure 13 in Kim et al. [19]. For our BCM and BCM-deep, we use ancestral sampling when NFE=2, zigzag sampling when NFE=3, and the combination of both when NFE=4. Our results indicate that ancestral and zigzag sampling can individually improve FID, and their combination can achieve even better performance.
@@
Figure 3: Uncurated samples generated by BCM-deep. We can observe that our proposed sampling methods can improve the FID while preserving the content to a large degree.
We first apply BCM for image generation on CIFAR-10 dataset [28]. We showcase some samples in Figure 3, and include more samples in Appendix G. We report the NFE/FID/IS along with several baselines in Table 2. First, we conclude that both ancestral sampling and zigzag sampling can improve the sample quality, while the combination of both yields the best performance.
Second, we can see that our proposed BCM shows clear advantages compared with earlier diffusion models. BCM achieves comparable or even better results within at least one order of magnitude fewer NFEs, compared with score SDE [7], DDPM [5] and EDM [15]. As for the methods boosting sampling speed through fast samplers or distillation techniques, including DDIM [6], DPM-solver-fast [26] and Progressive Distillation [27], BCM also achieves better results with even fewer NFEs.
Within the consistency model family, for 1-step sampling, iCT [18] surpasses BCM. However, as the number of function evaluations (NFEs) increases, following the sampling strategies detailed in Section 3.3, BCM starts to outperform. Additionally, a significant advantage of BCM’s multi-step sampling is its determinism, as illustrated in Figure 3. This contrasts with CMs, which introduces fresh noise at each iteration, thereby influencing the content of the image.
On the other hand, our model’s performance still falls short of CTM’s [19]. However, we note that 1) the training of CTM heavily relies on the adversarial loss, which may impact the stability; 2) CTM uses LPIPS [20] as the distance measure, which can have feature leakage and cause performance overestimation, as pointed by Kynkäänniemi et al. [29] and Song and Dhariwal [18].

### v4

@@ 3.2 Bidirectional Consistency Training
We now discuss the training of BCM, which we dub as Bidirectional Consistency Training (BCT). Following Song et al. (2023); Song & Dhariwal (2024), we discretize the time horizon $[0,T]$ into $N-1$ intervals with boundaries $0=t_{1}<t_{2}<\dots<t_{N}=T$.
@@
where ${\mathbf{x}}$ is one training sample, ${\mathbf{z}}\sim\mathcal{N}(0,{\bm{I}})$, and $\bar{\bm{\theta}}$ is the stop gradient operation on $\bm{\theta}$, and $\lambda(t_{n})=\nicefrac{{1}}{{|t_{n}-t_{n+1}|}}$. We replace $\bm{\theta}^{-}$ in Equation 3 with $\bar{\bm{\theta}}$ according to Song & Dhariwal (2024).
@@
However, Equation 7 will have different scales for different $u$ values, leading to a high variance during training. Therefore, inspired by Kim et al. (2024), we map both $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t},t,u)$ and ${\mathbf{x}}_{u}$ to time 00, and minimize the distances between these two back-mapped images, i.e.,
@@
where $\bar{\bm{\theta}}$ is the same $\bm{\theta}$ with stop gradient operation. We denote this as a “soft” trajectory constraint. Unfortunately, directly minimizing Equation 8 without a pretrained DM is still problematic. This is because, without a pretrained DM, we can only generate ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{u}$ from the diffusion SDE, i.e., by adding Gaussian noise to ${\mathbf{x}}$. However, ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{u}$ generated by the diffusion SDE do not necessarily lie on the same PF ODE trajectory, and hence Equation 8 still fails to build the desired bidirectional consistency. Instead, we notice that when the CT loss defined in Equation 6 converges, we have $\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}_{u},u,0)\approx\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}_{t},t,0)\approx{\mathbf{x}}$. We therefore optimize:
@@
Empirically, we found Equation 9 plays a crucial role in ensuring accurate inversion performance. We provide experimental evidence for this loss choice in Section B.2.1.
@@
$\begin{split}\mathcal{L}_{BCT}^{N}(\bm{\theta})=\mathbb{E}_{{\mathbf{z}},{\mathbf{x}},t_{n},t_{n^{\prime}}}\Big{[}\lambda(t_{n})&d(\bm{f}_{\bm{\theta}}({\mathbf{x}}+{t_{n+1}}{\mathbf{z}},t_{n+1},0),\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}+{t_{n}}{\mathbf{z}},t_{n},0))\\ +\lambda^{\prime}(t_{n},t_{n^{\prime}})&d\left(\bm{f}_{\bar{\bm{\theta}}}(\bm{f}_{\bm{\theta}}({\mathbf{x}}+t_{n}{\mathbf{z}},t_{n},t_{n^{\prime}}),t_{n^{\prime}},0)),\bm{f}_{\bar{\bm{\theta}}}({\mathbf{x}}+t_{n}{\mathbf{z}},t_{n},0)\right)\Big{]}\end{split}$ (10)
@@
Training from scratch. We can train BCM from scratch using Equation 10. We empirically found that all the training settings used for iCT (Song & Dhariwal, 2024), including the scheduler function for $N$, the sampling probability for $t_{n}$ (aka the noise schedule $p(n)$ in (Song et al., 2023; Song & Dhariwal, 2024)), the EMA rate, and more, also works well for BCM. Please refer to Section C.1 for more details on these settings and hyperparameters. We summarize the training process in Algorithm 1 and compare the training of CT, CTM, and BCT in Table 2 in Appendix E.
Bidirectional Consistency Fine-tuning. To reduce the training cost and increase scalability, we can also initialize BCM with a pretrained consistency model and fine-tune it using Equation 10 to incorporate bidirectional consistency. Recall that BCM takes in three arguments: 1) the sample ${\mathbf{x}}_{t}$ at time $t$, 2) the current time step $t$, and 3) the target time step $u$, and outputs the sample at time $u$, i.e., ${\mathbf{x}}_{u}$. In contrast, the standard consistency model only takes the first two inputs and can be viewed as a special case of BCM when the target time step $u$ is set to 00. Therefore, we aim to initialize BCM such that it preserves the performance of the pretrained CM when $u=0$. To achieve this, we concatenate the embeddings of $t$ and $u$, then pass the result through a linear layer (without bias vector) to reduce the dimensionality by half. The weight matrix is initialized as $[{\mathbf{I}},\bm{0}]$. The embedding layers for $t$ and $u$ are initialized using the pretrained CM’s embedding layer, but they are not tied together during fine-tuning. Other layers are initialized identically to those of the CM. It is easy to check that this initialization effectively ignores $u$, preserving the CM’s generation performance without fine-tuning. See Section C.1 for more details.
@@ 3.3 Sampling
In this section, we describe the sampling schemes of BCM. Similar to CMs and CTM, BCM supports 1-step sampling naturally. However, BCM’s capability to navigate both forward and backward along the PF ODE trajectory allows us to design more complicated multi-step sampling strategies to improve the sample quality. We present two schemes and a combined approach that has empirically demonstrated superior performance in the following.
@@
Ancestral Sampling. The most straightforward way for multi-step sampling is to remove noise sequentially. Specifically, we first divide the time horizon $[0,T]$ into $N$ sub-intervals with boundaries $0=t_{0}<t_{1}<\dots<t_{N}=T$. Then, we sample a noise image ${\mathbf{x}}_{T}\sim\mathcal{N}(0,T^{2}{\bm{I}})$, and sequentially remove noise with the network: ${\mathbf{x}}_{t_{n-1}}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t_{n}},t_{n},t_{n-1}),\quad n=N,N-1,...,1$. We note that the discretization strategy may differ from the one used during the training of BCM. Since this sampling procedure can be viewed as drawing samples from the conditional density $p_{{\mathbf{x}}_{t_{n-1}}|{\mathbf{x}}_{t_{n}}}({\mathbf{x}}_{t_{n-1}}|{\mathbf{x}}_{t_{n}})$, we dub it as ancestral sampling, and summarize it in Algorithm 2. We can also view 1-step sampling as ancestral sampling, where we only divide the time horizon into a single interval.
@@
Zigzag Sampling. Another effective sampling method (Algorithm 1 in Song et al. (2023)) is to iteratively re-add noise after denoising. Similar to ancestral sampling, we also define a sequence of time steps $t_{1}<\dots<t_{N}=T$. However, different from ancestral sampling where we gradually remove noise, we directly map ${\mathbf{x}}_{T}$ to ${\mathbf{x}}_{0}$ by $\bm{f}_{\bm{\theta}}({\mathbf{x}}_{T},T,0)$. We then add a fresh Gaussian noise to ${\mathbf{x}}_{0}$, mapping it from time 00 to time $t_{n-1}$, i.e., ${\mathbf{x}}_{t_{n-1}}={\mathbf{x}}_{0}+t_{n-1}\bm{\sigma}$, where $\bm{\sigma}\sim\mathcal{N}(0,{\bm{I}})$. This process repeats in this zigzag manner until all the designated time steps are covered. The two-step zigzag sampler effectively reduces FID in CMs (Song et al., 2023) and is theoretically supported (Lyu et al., 2023). However, the fresh noise can alter the content of the image after each iteration, which is undesirable, especially considering that our tasks will both include generation and inversion. One may immediately think that setting the injected noise to be the same as the initial random noise can fix this issue. However, we reveal that this significantly damages the quality of the generated images.
@@
Fortunately, our proposed BCM provides a direct solution, leveraging its capability to traverse both forward and backward along the PF ODE. Specifically, rather than manually reintroducing a large amount of fresh noise, we initially apply a small amount and let the network amplify it. In a nutshell, for iteration $n$ ($n=N,N-1,\cdots,2$), we have
${\mathbf{x}}_{0}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{t_{n}},t_{n},0),\quad{\mathbf{x}}_{\varepsilon_{n-1}}\leftarrow{\mathbf{x}}_{0}+\varepsilon_{n-1}\bm{\sigma},\quad{\mathbf{x}}_{t_{n-1}}\leftarrow\bm{f}_{\bm{\theta}}({\mathbf{x}}_{\varepsilon_{n-1}},\varepsilon_{n-1},t_{n-1}).$ (11)
@@
where $\varepsilon_{n-1}$ is the scale of the small noises we add in $n$-th iteration, and $\bm{\sigma}\sim\mathcal{N}(0,{\bm{I}})$ is a fresh Gaussian noise. We detail this scheme in Algorithm 3. To verify its effectiveness, in Figure 2, we illustrate some examples to compare the generated images by 1) manually adding fresh noise, 2) manually adding fixed noise, and 3) our proposed sampling process, i.e., adding a small noise and amplifying it with the network. We can clearly see our method maintains the generated content.
@@
Figure 2: Comparison of different strategies of adding fresh noise in zigzag sampling. (a) 1-step generation. (b) Zigzag sampling with manually added fresh noise, where the new noises drastically alter the content. (c) Zigzag sampling with manually added, fixed noise, i.e., we fix the injected fresh noise in each iteration to be the same as the initial one. We can see that the quality significantly deteriorates. (d) Zigzag sampling with BCM. At each iteration, we apply a small amount of noise and let the network amplify it. We can see that the image content is mostly maintained.
@@
Combination of Both. Long jumps along the PF ODE in zigzag sampling can lead to accumulative errors, especially at high noise levels, which potentially hampers further improvements in sample quality. Therefore, we propose a combination of ancestral sampling and zigzag sampling. Specifically, we first perform ancestral sampling to rapidly reduce the large initial noise to a more manageable noise scale and then apply zigzag sampling within this small noise level. We describe this combined process in Algorithm 4. We empirically found this combination results in superior sample quality compared to employing either ancestral sampling or zigzag sampling in isolation. We provide ablation to evidence the effectiveness of this combination in Section B.2.3.
@@ 3.4 Inversion
BCM inverts an image following the same principle of sampling. Specifically, we also set an increasing sequence of noise scales $\varepsilon=t_{1}<t_{2}<\cdots<t_{N}\leq T$. Note that, in contrast to the generation process, it is not always necessary for $t_{N}$ to equal $T$. Instead, we can adjust it as a hyperparameter based on the specific tasks for which we employ inversion. Then, given an image ${\mathbf{x}}_{0}$, we first inject a small Gaussian noise by ${\mathbf{x}}_{t_{1}}={\mathbf{x}}_{0}+\varepsilon\bm{\sigma}$, and then sequentially add noise with the network, i.e., ${\mathbf{x}}_{t_{n+1}}=f({\mathbf{x}}_{t_{n}},t_{n},t_{n+1}),\quad n=1,2,\cdots,N-1.$ The adoption of small initial noise is due to the observation that the endpoint of the time horizon is less effectively covered and learned during training, as discussed in Section B.2.4. Empirically, we find this minor noise does not change the image’s content and leads to lower reconstruction errors when $\varepsilon\approx 0.07$. One may also include denoising steps interleaved with noise magnifying steps, like zigzag sampling, but we find it helps little in improving inversion quality. We summarize the inversion procedure in Algorithm 5.
@@ 4 Experiments and Results
In this section, we first present the results of the two basic functions of BCM, i.e., generation and inversion. Then in Section 4.3, we show some potential applications enabled by these two functions.
@@ 4.1 Image Generation
Table 1: Sample quality on CIFAR-10 (left) and ImageNet-64 (right). We train BCM from scratch on CIFAR-10 and fine-tune it using our reproduced iCT model on ImageNet-64. Results estimated from Figure 13 in Kim et al. (2024). For our BCM and BCM-deep, we use ancestral sampling when NFE=2, zigzag sampling when NFE=3, and the combination of both when NFE=4. Our results indicate that ancestral and zigzag sampling can individually improve FID, and their combination can achieve even better performance. ∗∗Results by our reproduction.
@@
Diffusion with Fast Sampler / Distillation / Fine-tuning
DDIM (Song et al., 2021a) 10 8.23 -
DPM-solver-fast (Lu et al., 2022a) 10 4.70 -
AMED-plugin (Zhou et al., 2023) 5 6.61 -
Progressive Distillation (Salimans & Ho, 2022) 1 8.34 8.69
Diff-Instruct (Luo et al., 2024) 1 4.53 -
CD (LPIPS, Song et al., 2023) 1 3.55 9.48
@@
EDM (Karras et al., 2022) 35 2.04 9.84
CT (LPIPS, Song et al., 2023) 1 8.70 8.49
@@
CTM (LPIPS, GAN loss, Kim et al., 2024) 1 2.39 -
iCT (Song & Dhariwal, 2024) 1 2.83 9.54
@@
METHOD NFE ($\downarrow$) FID ($\downarrow$)
Diffusion with Fast Sampler / Distillation / Fine-tuning
DDIM (Song et al., 2021a) 10 18.3
DPM-solver-fast (Lu et al., 2022a) 10 7.93
AMED-solver (Zhou et al., 2023) 5 10.74
Progressive Distillation (Salimans & Ho, 2022) 1 7.88
Diff-Instruct (Luo et al., 2024) 1 5.57
EM Distillation (Xie et al., 2024) 1 2.20
CD (LPIPS, Song et al., 2023) 1 6.20
2 4.70
CTM (LPIPS, GAN loss, Kim et al., 2024) 1 1.92
2 1.73
ECT-XL (Geng et al., 2024) 2 1.67
Direct Generation
EDM2-L/XL (Karras et al., 2024) 63 1.33
CT (LPIPS, Song et al., 2023) 1 13.0
2 11.1
iCT (Song & Dhariwal, 2024) 1 4.02 (4.60∗∗)
2 3.20 (3.40∗∗)
iCT-deep (Song & Dhariwal, 2024) 1 3.25 (3.94∗∗)
2 2.77 (3.14∗∗)
BCM (ours) 1 4.18
2 2.88
3 2.78
4 2.68
BCM-deep (ours) 1 3.14
2 2.45
3 2.61
4 2.35
@@
Figure 3: Samples by BCM-deep on CIFAR-10.
            (a) 1-step (FID 3.14)
            (d) 4-step (FID 2.35)
Figure 4: Samples by BCM-deep on ImageNet-64.
We first evaluate the image generation performance of BCM. On CIFAR-10, we train our BCM from scratch, while on ImageNet-64, we first train an iCT model, and fine-tune BCM from it, as we discussed in Section 3.2. When reproducing iCT on ImageNet-64, we encountered instability during training, and found it difficult to reproduce the reported FID by Song & Dhariwal (2024). To mitigate the performance gap, we removed the attention layer at the 32x32 resolution and added normalization to the QKV attention, following (Karras et al., 2024). We empirically find these simple modifications bring positive influences on performance. We include the hyperparameters for training and sampling in Sections C.1 and C.2. We report image quality and NFE for CIFAR-10 and ImageNet-64 in Table 1, and we visualize some generated samples Figures 3, 4 and H. As we can see,
  • both ancestral sampling and zigzag sampling can improve the sample quality, while the combination of both further yields the best performance;
  • our proposed BCM achieves comparable or even better results within at least one order of magnitude fewer NFEs compared with diffusion models; BCM presents better results with even fewer NFEs then the methods boosting sampling speed through fast samplers; it is also comparable to modern distillation approaches like EM distillation by Xie et al. (2024);
  • comparing with CMs, iCT (Song & Dhariwal, 2024) surpasses BCM for 1-step sampling. This is not surprising, given that BCMs tackle a significantly more complex task compared to CMs. However, as the number of function evaluations (NFEs) increases, following the sampling strategies detailed in Section 3.3, BCM starts to outperform. This indicates that our approach allows the model to learn bidirectional consistency without hurting the generation performance;
  • our model’s performance still falls short of CTM’s (Kim et al., 2024). However, CTM relies heavily on adversarial loss and uses LPIPS (Zhang et al., 2018) as the distance measure, which can have feature leakage (Song & Dhariwal, 2024; Kynkäänniemi et al., 2023). ECT (Geng et al., 2024) also outperforms BCM on ImageNet-64 with 2 NFEs. However, we note that their approach—fine-tuning from a pretrained diffusion and using the advanced EDM2 architecture (Karras et al., 2024)—is orthogonal to ours. Our method could be combined with theirs for potential improvements in both performance and training speed.
Another interesting phenomenon we observed on ImageNet-64 is that our fine-tuned BCM actually outperforms its initialization (i.e., our pretrained iCT) on generation. In contrast, we empirically find fine-tuning iCT with the same training budget did not yield any improvement. We suspect this is due to the “soft” trajectory constraint in the bidirectional training objective, which may act as a form of regularization, aiding optimization. We leave further investigation of this for future work.